iKAGRA ETM suspension preparation log
- 12th Jan- (2 weeks) Disassembling, Washing, and transportation
- 25th Jan - (2 weeks) Assembling and Test Hanging with an Al dummy mirror
- 7th Feb- (1 week) coil winding, measurement of transfer function, transportation to the End area and installation test
15th Feb- (1 week)installation
Suspensions and mirrors were installed on 17th Feb so this 6 weeks task was completed on schedule.
2. Disassembling in the ATC clean booth at Mitaka (12th-19th Jan.)
- One suspensions was moved from the desiccator at the west end of TAMA and the other was moved from the vacuum tank.
- Two nuts stuck. Hirata-san removed them.
- A Cu screw stuck. Hirata-san removed it.
3. Washing and transportation to Kamioka
The ultra sonic washer at ATC was broken on 20th Jan.
A new washer was ordered on 22th(Fri) and arrived at Mitaka on 25th(Mon).
Eight screws for adjusting the IM holding plate was dirty.
Takahashi-san ordered similar screws.
4. New parts preparation
4.1 Coil bobbin (Delivered to Kamioka on 4th Feb.)
- The outer diameter of the coil has been changed from 14.2mm to 14.5mm.
- Due to this change, the number of winding becomes 39 from 41.
- Wind 13 times / 1 layer.
4.2 Coil winding (8th-16th Feb.)
[Kuwahara-san, and Ushiba-san from U-Tokyo, Lee-san from Okazaki, and Ohishi]
- We used coil winding machine.
- Lead wires from the bobbin to Dsub connector were too short (20cm). The length must be 40cm or more.
- At the end of the lead wires, D-sub pins are swaged. We have run short of pins and waiting for delivery on 29th Feb(Mon).
- Peeling off the coating of wires longer than the length of those pins make the wire broken easily.
4.3 D-sub connector plates (8th Feb.)
- M3 holes were changed to phi3.1 holes. Yamamoto-san used drilling machine at Kamioka surface office.
- One must be careful about the pillar position in order to attach this plate to the suspension.
4.4 15pin-9pin changer for coil driver output [Ushiba-san, Ohishi]
4.5 Screws in replace with Cu screws [Ohishi]
- Replaced all Cu screws to stainless screws. Magnetism problem was indicated.
4.6 D-sub 9pin - BNC changer [Doi-san]
- Used to input the signal from the laser displacement measurement system to AA/ADC of the digital system.
5.1 Magnet and standoff fixture jig
- We used
5.2 Dock (mirror hanging jig)
(4 ; with the diameter of 2mm and the length of 5mm NdFeB magnets) [Tomaru-san, Yamamoto-san]
6. Some other items
6.1 Pico-motors and PZTs
- We have 4 pico-motors except for x-axis (the direction of the light ) y, z, pitch, yaw
6.2 Coil drivers
- We use KAGRA coil driver; we need to cheek its performance.
- Two TAMA PRM mirrors (with about 50% and 90% reflectivities each, resultant contrast becomes 96%) were delivered to Kamioka on 12th Feb.
6.5 Mirror container [Tatsumi-san]
- Some plates once used for transportation of MC mirrors were used
- Two places of the container were cut to avoid standoffs.
- Vacuum compatibility?
[Tomaru-san, Ohishi, Yamamoto-san]
- On 21th, suspension frame, security frame, xy stages completed
[Yamamoto-san, Uchiyama-san, Doi-san, and Ohishi]
- Hanging jig preparation
9. Transfer Function Measurement (8th-16th Feb.)
[Shoda-san, Hayama-san, Ushiba-san, Kuwahara-san, Kokeyama-san, Kuwahara-san, Tomaru-san, Lee-san, and Ohishi]
- First Shoda-san kindly made a simple model under PR2 digital system for this test.
10. Preparation of End area and suspension transportation to the End
[Kimura-san, Sawada-san, Tomaru-san, Prof. Suzuki, Telada-san, Uchiyam-san]
- Uchiyama-san bought many things we need at the end area and KEK member transported them and cleaned the work area.
11. Installation [Akutsu-san, Uchiyama-san, Yamamoto-san, Miyoki-san, and Ohishi ] (15th-17th Feb.)
11.1 Suspension installation
- Suspension was installed by Miyoki-san, Akutsu-san and Yamamoto-san on 15th Feb at the Yend.
- Suspension and mirrors were installed by at the Xend.